OpenStudy (anonymous):

integrate (cos2x)/(cosx+sinx)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ \cos^2(x)-\sin^2(x) }{ \cos(x)+\sin(x) }=\int\limits_{}^{}\cos(x)-\sin(x)= \sin(x)+\cos(x)+c\]